1. Calculate the pH of water in a cloud that has equilibrated with SO2(g) at a concentration of 1.0 ppm by volume.
  2. A solution contains 2 mg/L CN- and is at pH 8.5
  • What is the concentration of HCN(aq) in solution?
  • If the partial pressure of HCN(g) above the solution is 10-6 atm, would HCN transfer from the gas phase into solution or from the solution into the gas, or is the system at equilibrium?
1
Expert's answer
2021-10-22T03:51:59-0400

PSO2 = (1 atm)(1x 10-6) = 1 x 10-6 atm

[H2SO3] = KH·PSO2 = (1.0 M/atm)(1 x 10-6 atm) = 1 x 10-6M

Reaction: H2SO3 ↔ H++ HSO3-



Ka1 = 0.017

Note: although we can not really use an ICE (initial change equilibrium) table because

SO2 will not be depleted from the gas phase and H2SO3 will be constant, we can expect

that [H+] = [HSO3-] (assuming no other sources of H+


)

Ka1 = 0.017 = [H+][HSO3-]/[H2SO3] = [H+]2/1x 10-6



[H+] = [(0.017)(1 x 10-6)]0.5 = 1.84 x 10-4 M

pH = 3.73